nikkixmarie:

A car travels 1/6 of the distance between two cities in 3/5 of an hour. At this rate, what fraction of the distance between the two cities can the car travel in 1 hour?

nikkixmarie:

@Hero

Hero:

@nikkixmarie how much is 3/5 of an hour?

nikkixmarie:

36 mins

Hero:

Okay, so this is a proportion. Do you know what that is?

nikkixmarie:

yes

Hero:

Would you like to try setting it up?

nikkixmarie:

idk how to start it

Hero:

\(\dfrac{1/6}{x} = \dfrac{36}{60}\)

nikkixmarie:

so do i divide 36/60

Hero:

Yes, 36/60 can be reduced.

nikkixmarie:

3/5

Hero:

Okay so now you have \(\dfrac{1/6}{x} = \dfrac{3}{5}\)

Hero:

Now what do we do?

nikkixmarie:

find x i think?

Hero:

Correct.

nikkixmarie:

idk how...

nikkixmarie:

wait no 5/18

Hero:

Good. That's the correct answer.

nikkixmarie:

ok thank you